OpenStudy (anonymous):

show your work please i do not get it at "ALL"! x2 - 3|x - 2| - 4x = - 6

myininaya (myininaya):

\[\text{if } x-2>0 \text{ then} |x-2|=x-2\] \[\text{if } x-2<0 \text{ then} |x-2|=-(x-2)\]

myininaya (myininaya):

so we have two cases

myininaya (myininaya):

solve: \[x^2-3(x-2)-4x=-6\] \[x^2-3(-(x-2))-4x=-6\]

myininaya (myininaya):

you have two cases so two equations to solve